# Physics-informed Hamiltonian learning for large-scale optoelectronic property prediction

### v1

@@ Discussion
While in this work we deliberately adopted a simple ML architecture, future extensions could tackle physics-informed frameworks that leverage more powerful global approaches, such as ENNs, to capture non-local electronic effects beyond the present description. Similarly, although our pilot implementation presented in this work relied on semi-local DFT, the modest requirement for first-principles input data makes it feasible to incorporate more advanced electronic-structure methods, including more accurate hybrid functionals or many-body perturbation theory, which could further enhance prediction accuracy. These directions, together with applications to broader classes of complex materials, highlight the potential of physics-informed Hamiltonian learning as a versatile tool for advancing quantitative materials modeling under increasingly realistic conditions.
@@ Methods
For CsPbBr3, we employed a $2\times 2\times 2$ supercell, $3\times 3\times 3$ $k$-point grid, Tkatchenko–Scheffler van der Waals corrections [48], and a time step of $8\text{\,}\mathrm{f}\mathrm{s}$ to perform a DFT-MD production run containing 2,000 steps at each temperature. For large-scale MD calculations of CsPbBr3, we trained an MLFF in VASP [5, 49, 50] at $425\text{\,}\mathrm{K}$ using the Bayesian on-the-fly scheme and a time step of $6\text{\,}\mathrm{f}\mathrm{s}$ for a $4\times 4\times 4$ supercell of the cubic structure. In order to ensure robustness and convergence of the resulting MLFFs, the training runs lasted $100\text{\,}\mathrm{p}\mathrm{s}$. During training, the default set of hyperparameters was first employed, specifying the following tags in Vasp: ML_RCUT1 = $8\text{\,}\AA$ , ML_RCUT2 = $5\text{\,}\AA$ , ML_SION1 = $0.5\text{\,}\frac{}{}$, ML_SION2 = $0.5\text{\,}\frac{}{}$. Subsequently, the atomic local reference configurations were refined using ML_MODE = SELECT in Vasp together with an optimized set of hyperparameters, specified as follows: ML_RCUT1 = $12\text{\,}\AA$ , ML_RCUT2 = $10\text{\,}\AA$ , ML_SION1 = $0.5\text{\,}\frac{}{}$, ML_SION2 = $0.5\text{\,}\frac{}{}$. The final MLFFs were then used to generate trajectories for the $16\times 16\times 16$ supercell of CsPbBr3 at a temperature of $425\text{\,}\mathrm{K}$.
@@
where $\gamma$ is a learnable parameter that depends on the atomic species, and $\hat{L}$ and $\hat{S}$ denote the orbital and spin-angular momentum operators, respectively. In the case of a basis of AOs given in spherical harmonics, we can analytically evaluate Eq. 6 (see, e.g., Ref. 53). For systems described using hybrid orbitals (HOs), as employed in our previous work on GaAs [26], the SOC Hamiltonian in the HO basis is computed via the transformation
@@
We separately trained and optimized each of the three models, TB, ML, and SOC. Each optimization consisted of two steps: a hyperparameter optimization (if applicable) using Tree-structured Parzen Estimator [54] followed by a parameter optimization using gradient descent with the ADAM optimizer [55].
For each material, the TB model was optimized using only the energy eigenvalue data for the unit cell of the system. We first optimized the effective core charge parameter (see Sec. C in Ref. 26) in the hyperparameter optimization using 200 optimization and 300 hyperoptimization steps. We used learning rates between 0.1 and 0.3 and initialized all TB parameters as 1. Next, we obtained optimal TB parameters in a longer optimization run. These final TB parameters were kept constant in all subsequent calculations of each material.
The ML model was optimized using energy eigenvalue data for selected snapshots from MD trajectories at a single temperature as described above. From the entire set of descriptor vectors for all training structures, a subset of $N_{\mathrm{p}}$ kernel support points is sampled using k-means clustering with $N_{\mathrm{cluster}}$ clusters, in which both $N_{\mathrm{p}}$ and $N_{\mathrm{cluster}}$ act as hyperparameters. To account for the varying importance of each cluster, we weighted the number of points sampled from the $i$-th cluster by
$w_{i}=\alpha s_{i}+(1-\alpha)v_{i},$ (8)
where $s_{i}$, $v_{i}$ are the normalized size and variance of the $i$-th cluster, respectively, and $\alpha=0.5$ for all our calculations. At least one point was sampled from each cluster to ensure that each type of interaction was sampled at least once. The ML parameters were initialized as 0, and we used a learning rate of 0.01. After hyperparameter optimization, we performed one training run of 200 optimization steps where SOC and ML parameters were fitted together. Afterward, we reinitialized the ML model and refitted its parameters, keeping SOC parameters fixed, in an additional 200 steps. For CsPbBr3 and GaAs, structural entries in the descriptor were kept fixed. For MAPbBr3, they were updated dynamically to accommodate thermal changes of the lattice with temperature. For this, the dynamic structural entries in the descriptor were normalized by the cutoff radius or by the maximum angle $2\pi$, and the environment as defined in Eq. 5 was allowed to scale by a fixed factor, env_scale, serving as another hyperparameter. The final ML models comprised 250 parameters for GaAs, 640 for CsPbBr3, and 750 for MAPbBr3.
To perform gradient-based optimization, the gradient of the loss function with respect to each model parameter is required, which can be obtained via the chain rule. Specifically, we are interested in computing
@@
where $p$ denotes a model parameter and R are lattice translation vectors.
@@
$L=\frac{1}{\Sigma_{w_{n}}\Sigma_{w_{\textbf{k}}}}\sum_{n\textbf{k}}w_{n}w_{\textbf{k}}|E^{\mathrm{target}}_{n\textbf{k}}-E^{\mathrm{pred}}_{n\textbf{k}}|,$ (10)
@@
As described in detail above, the training and validation datasets consisted of energy eigenvalues obtained from a variety of atomic configurations, which may come from MD trajectories or simply from the primitive unit cell. As also described above, we evaluated eigenvalues at a set of k{k}-points distributed across the irreducible Brillouin zone, using a $\Gamma$-centered grid. The choice of the energy window containing the bands used in the optimization must be made carefully, as it depends on the pseudopotential used in the DFT calculations. For CsPbBr3 and MAPbBr3, we included the nine uppermost valence bands together with additional conduction bands to account for contributions of the remaining orbitals (four for CsPbBr3, three for MAPbBr3). When computing the loss, we assigned increased weight to the $\Gamma$-point as well as to the highest-energy valence and lowest-energy conduction bands, while reducing the weight of the two uppermost bands. These numbers were doubled when SOC is included in the calculation and further scaled with the size of the supercell. The reason for this is that these bands typically exhibit larger errors, which can disproportionately influence the fit as the model attempts to minimize their contributions. Moreover, they may include components from higher-lying conduction bands due to band entanglement, making them difficult for the model to predict with sufficient accuracy. Since these states lie far from the band gap region, they are also less critical for most target applications, further justifying their reduced influence during training.
@@
The datasets generated during the current study are available from the corresponding author upon reasonable request.
@@
The code used in this study is available at GitHub (https://github.com/TheoFEM-TUM/Hamster.jl)
@@ Acknowledgments
We thank Kaiwen Chen (TU Munich) for fruitful discussions. Funding provided by the Bavarian Collaborative Research Project Solar Technologies Go Hybrid (SolTech), the Deutsche Forschungsgemeinschaft (DFG) via SPP2196 Priority Program (Project-ID: 424709454), Germany’s Excellence Strategy-EXC 2089/1-390776260, and the Studienstiftung des Deutschen Volkes, are gratefully acknowledged. The authors further acknowledge the Gauss Centre for Supercomputing e.V. for funding this project by providing computing time through the John von Neumann Institute for Computing on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre.

### v2

@@ Discussion
An important consideration that follows from this capability is chemical transferability. In principle, transfer learning is feasible when the source and target systems share similar local environments and orbital character, allowing pre-trained parameters to serve as meaningful initializations. This may hold for materials that differ only by modest compositional changes or chemically similar substitutions. However, when electronic structure or orbital symmetries differ substantially, previously learned parameters are unlikely to remain valid, and retraining becomes necessary. At the same time, these considerations highlight the longer-term potential for developing a more general model that spans broader classes of materials and can be efficiently adapted to new systems with minimal additional training.
While in this work we deliberately adopted a simple ML architecture, future extensions could tackle physics-informed frameworks that leverage more powerful global approaches, such as ENNs, to capture non-local electronic effects beyond the present description. Similarly, although our pilot implementation presented in this work relied on semi-local DFT, the modest requirement for first-principles input data makes it feasible to incorporate more advanced electronic-structure methods. These could include hybrid functionals or many-body perturbation theory, potentially extending the framework to including excited-state effects that can play an important role in the modeling of optoelectronic properties. Such enhancements could further enhance prediction accuracy. These directions, together with applications to broader classes of complex materials, highlight the potential of physics-informed Hamiltonian learning as a versatile tool for advancing quantitative materials modeling under increasingly realistic conditions.
@@ Methods
For CsPbBr3, we employed a $2\times 2\times 2$ supercell, $3\times 3\times 3$ $k$-point grid, Tkatchenko–Scheffler van der Waals corrections [50], and a time step of $8\text{\,}\mathrm{f}\mathrm{s}$ to perform a DFT-MD production run containing 2,000 steps at each temperature.
For large-scale MD calculations of CsPbBr3, we first trained three MLFFs in VASP [5, 51, 52], at $425\text{\,}\mathrm{K}$, $525\text{\,}\mathrm{K}$, and $625\text{\,}\mathrm{K}$, respectively, using the Bayesian on-the-fly scheme and a time step of $6\text{\,}\mathrm{f}\mathrm{s}$ for a $4\times 4\times 4$ supercell of the cubic structure. In order to ensure robustness and convergence of the resulting MLFFs, the training runs lasted $100\text{\,}\mathrm{p}\mathrm{s}$. During training, the default set of hyperparameters was first employed, specifying the following tags in Vasp: ML_RCUT1 = $8\text{\,}\AA$ , ML_RCUT2 = $5\text{\,}\AA$ , ML_SION1 = $0.5\text{\,}\frac{}{}$, ML_SION2 = $0.5\text{\,}\frac{}{}$. Using the DFT data collected during on-the-fly training, we trained three MACE models [9], one per temperature, using 2 message passing layers, messages with 128 channels and max_L of 2, correlation order of 3, spherical harmonics up to $l=3$, and a radial cutoff of $10\text{\,}\AA$ . The weighted mean squared error loss function was implemented with initial energy and force weights being 1 and 100, respectively, increasing to 1000 and 100 in the last 20% of the epochs, with stochastic weight averaging enabled. The MACE models were trained with a batch size of 5 for a total of 600 epochs. The validation sets were chosen to be 10% of the configurations in each data set. The reference energies for isolated atoms were obtained through spin-polarized DFT calculations. The resulting MACE force fields were used with the Lammps code [53], and applied to a $16\times 16\times 16$ supercell of the cubic structure for 5000 time steps of 0.6 fs.
@@
where $\gamma$ is a learnable parameter that depends on the atomic species, and $\hat{L}$ and $\hat{S}$ denote the orbital and spin-angular momentum operators, respectively. In the case of a basis of AOs given in spherical harmonics, we can analytically evaluate Eq. 7 (see, e.g., Ref. 55). For systems described using hybrid orbitals (HOs), as employed in our previous work on GaAs [26], the SOC Hamiltonian in the HO basis is computed via the transformation
@@
Sampling of kernel support points
From the entire set of descriptor vectors for all training structures, a subset of $N_{\mathrm{p}}$ kernel support points is sampled using k-means clustering with $N_{\mathrm{cluster}}$ clusters. To account for the varying importance of each cluster, we weighted the number of points sampled from the $i$-th cluster by
$w_{i}=\alpha s_{i}+(1-\alpha)v_{i},$ (9)
where $s_{i}$, $v_{i}$ are the normalized size and variance of the $i$-th cluster, respectively, and $\alpha=0.5$ for all our calculations. At least one point was sampled from each cluster to ensure that each type of interaction was sampled at least once.
@@
We perform separate training runs to optimize each set of model parameters, thereby avoiding numerical instability. Specifically, we optimize $V_{v}$ for the TB model (Eq. 3), $h_{n}$ for the ML model (Eq. 4), and the $\gamma$ parameters for the SOC model (Eq. 7) using gradient descent with the ADAM optimizer [56]. For the TB parameters, learning rates between 0.1 and 0.3 were found to work well, and all TB parameters were initialized to 1. The ML corrections are expected to be small, so a smaller learning rate of 0.01 or 0.005 is used, with all ML parameters initialized to 0. SOC parameters are also initialized to 1. Since the SOC model has only a few parameters, it is largely insensitive to the learning rate and converges within just a few iterations.
We optimize the TB model using the non-distorted primitive unit cell and using 12 randomly selected snapshots from MD as validation data to prevent overfitting; the ML and SOC models are optimized using 12 randomly selected snapshots from an MD run at a single temperature, as described above.
The TB model is optimized for up to 1000 training steps, with early stopping to prevent overfitting. For the ML and SOC models, we first run a combined training of 200 optimization steps, during which both ML and SOC parameters are fitted simultaneously. Next, the ML parameters are reinitialized to 0 and refitted for an additional 200 steps while keeping the SOC parameters fixed. This two-stage procedure is practical because the ML and SOC models strongly influence the band gap and are interdependent.
At each iteration, the gradient of the loss function with respect to each parameter is required. Using the chain rule, this can be expressed as
@@
where $p$ denotes a model parameter and R are the lattice translation vectors.
@@
$L=\frac{1}{\sum_{n}{w_{n}}\sum_{\textbf{k}}{w_{\textbf{k}}}}\sum_{n\textbf{k}}w_{n}w_{\textbf{k}}|E^{\mathrm{target}}_{n\textbf{k}}-E^{\mathrm{pred}}_{n\textbf{k}}|,$ (11)
@@
Nevertheless, our optimization strategy has some limitations. First, the TB, ML, and SOC parameters are trained in separate stages to avoid numerical instability, which prevents fully capturing their mutual interdependence. Consequently, subtle correlations between these components may not be entirely reflected in the final model. Second, gradient-based optimization may be less effective if the Hamiltonian exhibits strong nonlinear dependence on its parameters, and it cannot guarantee convergence to the global minimum. Despite these limitations, our approach provides a stable and computationally efficient framework for fitting effective Hamiltonians.
Hyperparameter Optimization
For parameters that cannot be efficiently optimized using gradient descent, we perform hyperparameter optimization using a Tree-structured Parzen Estimator (TPE) [57]. In each hyperparameter iteration, the TPE proposes a new candidate using a Bayesian estimator. For each candidate, a full parameter optimization is performed, with the model parameters reinitialized at the start of the run. Once the optimal hyperparameter setting is identified, a final optimization is carried out to obtain the corresponding optimal model parameters.
For the TB model, the effective core charge parameter (see Sec. C in Ref. 26) is treated as a hyperparameter, since computing its gradient is impractical. For the ML model, the hyperparameters include the number of kernel support points $N_{\mathrm{p}}$, the number of clusters $N_{\mathrm{cluster}}$, and the kernel width $\sigma$. As described in the main text, the structural entries in the descriptor can either be fixed or dynamically updated based on lattice distortions. For CsPbBr3 and GaAs, the structural entries are kept fixed. For MAPbBr3, they are updated dynamically to account for thermal lattice changes. In this case, the dynamic entries are normalized by the cutoff radius or the maximum angle $2\pi$, and the environment defined in Eq. 5 is scaled by a fixed factor, env_scale, which is treated as an additional hyperparameter. The final ML models comprise 250 parameters for GaAs, 910 for CsPbBr3, and 420 for MAPbBr3.
Band window selection and weighting
As described in detail above, the training and validation datasets consisted of energy eigenvalues obtained from DFT. Here, the choice of the energy window containing the bands used in the optimization must be made carefully, as it depends on the pseudopotential used in the DFT calculations and the number of Kohn-Sham orbitals is generally larger than the number of basis states in the TB basis. For CsPbBr3 and MAPbBr3, we included the nine uppermost valence bands together with additional conduction bands to account for contributions of the remaining orbitals (five for CsPbBr3, four for MAPbBr3). When computing the loss, increased weight was assigned to the $\Gamma$-point and to the highest valence and lowest conduction bands, while reducing the weight of the two uppermost bands. These numbers were doubled when SOC is included in the calculation and further scaled with the size of the supercell. The reason for this is that these bands typically exhibit larger errors, which can disproportionately influence the fit as the model attempts to minimize their contributions. Moreover, they may include components from higher-lying conduction bands due to band entanglement, making them difficult for the model to predict with sufficient accuracy. Since these states lie far from the band gap region, they are also less critical for most target applications, further justifying their reduced influence during training.
@@
The source data that support the findings of this study are available in the Zenodo repository with the identifier https://doi.org/10.5281/zenodo.18485403 [58].
@@
The code used in this study is available at GitHub (https://github.com/TheoFEM-TUM/Hamster.jl). The calculations were performed using version 0.2.1.
@@ Acknowledgments
We thank Kaiwen Chen and Jonas Oldenstaedt (TU Munich) for assistance with numerical aspects of the calculations. Funding provided by the Bavarian Collaborative Research Project Solar Technologies Go Hybrid (SolTech), the Deutsche Forschungsgemeinschaft via SPP2196 Priority Program (Project-ID: 424709454), Germany’s Excellence Strategy-EXC 2089/1-390776260, and the Studienstiftung des Deutschen Volkes, are gratefully acknowledged. The authors further acknowledge the Gauss Centre for Supercomputing e.V. for providing computing time through the John von Neumann Institute for Computing on the GCS Supercomputer JUWELS at Jülich Supercomputing Centre.
@@
Competing interests
The authors declare no competing interests.